the sum and difrence of two numbers are 32 and 16 respectively. find the numbers?

Answer by drpipespropmzy(2)

Let the numbers be x and y
Sum of numbers , x+y=32 equation 1
Difference of numbers , x-y=16 equation 2
Solving equations 1 and 2 simultaneously
Equation 1 minus equation 2

x+y=32
x-y=16
2y=16
y=8.
Putting y=8 into equations 1
x+8=32
x=32-8
x=24
Hence the numbers are 24 and 8